# Scintillating properties of Cs2ZrCl6 crystals in the temperature range of 5–300 K

### v2

@@ 1 Introduction
In the framework of the recently executed CHARM (radiopure Crystal with HAfnium for Rare Mode of alpha decay searches) initiative at Queen’s University (Kingston, Canada), Cs2ZrCl6 (CZC) crystals with dimensions of 2260 mm and mass of about 60 g were developed [10]. One of the main aim of this R&D was to develop an appropriate detector material to study rare nuclear decays that may occur in natural Zr isotopes, such as the neutrinoless 2 decay of 94,96Zr and strongly forbidden decay of 96Zr. The CZC compound is the first room-temperature ambient-tolerant crystal scintillator containing Zr in a relatively-high mass-fraction (16%). It allows to implement the “source = detector" approach, when the decaying isotope is embedded in the detector material, that leads to a high detection efficiency of the processes of interest. Moreover, these detectors exhibit a high identification capability for and particles, along with a reasonable energy resolution (Full Width at Half Maximum) FWHM = 5% at 662 keV line of a 137Cs source. Within long-term low-background measurements, novel experimental limits on the half-lives of various modes of the 2 decays of 96Zr and 94Zr were set at the level of T1/2 1017 – 1020 years [10]. This was the first complex study of CZC scintillating crystals in terms of their chemical purity and internal radioactive contamination, detailed analyses of their internal background components and scintillating performance at RT that all together confirm the potential of these crystals as a detector for rare decays of Zr isotopes.
@@ 2.1 Crystal production
The single crystal sample used in this study was obtained from Cs2ZrCl6 ingot (2260 mm, 60 g) produced through a multi-stage crystallization process by the vertical Bridgman-Stockbarger technique adopted from [10]. In this method, CsCl (99.9) and ZrCl4 (99.9) powders were used as starting materials. The latter was subjected to a two-stage sublimation process prior to the Cs2ZrCl6 compound synthesis. A pulling rate of 1.5 mm/hour and a temperature gradient of 25∘C/cm at the crystallization zone were maintained during the first crystal growth. Then, the obtained ingot was processed to remove all inclusions, as well as the first- and last-to-freeze sections. The second growth was performed with a 0.5 mm/hour pulling rate at the same temperature gradient. The final CZC single crystalline ingot was cut by a diamond wire saw in the way to obtain a non-oriented rectangular sample - 20146 mm, which then was processed with a 1200 grit sandpaper and mineral oil as a lubricant. Before installation, the sample was wrapped with four layers of a Teflon tape (except of the side facing the PMT and area facing the source) serving as a diffuse light reflector and as additional thermal insulation to prevent cracking of the crystal due to any possible sudden changes in temperature of the sample holder. Then it was secured in the sample holder using a few additional pieces of a Teflon tape.
@@ 3.1 Light yield over 5–300 K temperature interval
The mean of the fit provided a relative measurement of the light yield. The peak position of the ’s was found to be 30,650250 ADUs with a standard deviation of 1800300 ADUs while ’s peak position was determined to be 65,400100 ADUs with a standard deviation of 3,24040 ADUs. The uncertainties provided for light yield measurements includes only statistical uncertainties associated with the best fit parameters.
@@ 3.2 Pulse-shape analysis
Initially, events with a noisy baseline, pileups, and badly determined staring point were discarded from the collected dataset. Then, the different peak position on the detector energy scale for events induced by and irradiation was used to preliminary distinguish and select events of each type. See, for instance, the corresponding peak positions on the energy spectrum shown in figure 3. Events within two sigma around mean value of the corresponding peaks were selected for the further analysis. Only events with energy corresponding to a 662 keV full absorption peak were further considered as events in the pulse-shape analysis. Next, the mean time and prompt-to-total collected charge ratio were calculated for each event, and were used as additional discrimination and selection parameters to improve quality of the selected "pure" event of each type. These methods are well described in [31], where authors adopted those techniques to discriminate signals induced by particles from those of quanta in measurements with a CHC crystal.
Once "pure" events of each irradiation type were identified, the corresponding average waveforms were calculated. Then, the average waveform was fitted with a triple exponential function described by eq. 3.1, where , as previously shown in [13]. To fit the average ’s waveform was used a double exponential function provided by eq. 3.2, where . In both eq. 3.1 and eq. 3.2, represents the time duration of the average pulse, is the decay time-constant, and is the normalized relative intensity. The best fit of both types average waveforms normalized to their area, see figure 5, indicates that the average ’s waveform lasts longer than the average ’s waveform.
@@ 4 Discussion and conclusion
The intensity of the scintillating light emission under particles irradiation is rising when the CZC crystal is cooling down from RT to about 135 K reaching 19,7001,700 photons/MeV, and then the amount of the emitted light decreases to about 10,600900 photons/MeV at 75 K. While in the temperature interval from 75 K down to 5 K, it remains constant within uncertainties. It should be emphasized, that the present measurements confirms unusual behavior of the scintillating light intensity observed with the CZC crystal in [13], a decrease in the luminescence intensity as the temperature decreases from 125 K to 80 K, referred there as a "negative quenching".

### v3

@@ 1 Introduction
In the framework of the recently executed CHARM (radiopure Crystal with HAfnium for Rare Mode of alpha decay searches) initiative at Queen’s University (Kingston, Canada), Cs2ZrCl6 (CZC) crystals with dimensions of 22 60 mm and mass of about 60 g were developed [10]. One of the main aim of this R&D was to develop an appropriate detector material to study rare nuclear decays that may occur in natural Zr isotopes, such as the neutrinoless 2 decay of 94,96Zr and strongly forbidden decay of 96Zr. The CZC compound is the first room-temperature ambient-tolerant crystal scintillator containing Zr in a relatively-high mass-fraction (16%). It allows to implement the “source = detector” approach, when the decaying isotope is embedded in the detector material, that leads to a high detection efficiency of the processes of interest. Moreover, these detectors exhibit a high identification capability for and particles, along with a reasonable energy resolution (Full Width at Half Maximum) FWHM = 5% at 662 keV line of a 137Cs source. Within long-term low-background measurements, novel experimental limits on the half-lives of various modes of the 2 decays of 96Zr and 94Zr were set at the level of T1/2 1017–1020 years [10]. This was the first complex study of CZC scintillating crystals in terms of their chemical purity and internal radioactive contamination, detailed analyses of their internal background components and scintillating performance at RT that all together confirm the potential of these crystals as a detector for rare decays of Zr isotopes.
@@ 2.1 Crystal production
The single crystal sample used in this study was obtained from Cs2ZrCl6 ingot (22 60 mm, 60 g) produced through a multi-stage crystallization process by the vertical Bridgman-Stockbarger technique adopted from [10]. In this method, CsCl (99.9%) and ZrCl4 (99.9%) powders were used as starting materials. The latter was subjected to a two-stage sublimation process prior to the Cs2ZrCl6 compound synthesis. A pulling rate of 1.5 mm/hour and a temperature gradient of 25∘C/cm at the crystallization zone were maintained during the first crystal growth. Then, the obtained ingot was processed to remove all inclusions, as well as the first- and last-to-freeze sections. The second growth was performed with a 0.5 mm/hour pulling rate at the same temperature gradient. The final CZC single crystalline ingot was cut by a diamond wire saw in the way to obtain a non-oriented rectangular sample - 20 14 6 mm, which then was processed with a 1200 grit sandpaper and mineral oil as a lubricant. Before installation, the sample was wrapped with four layers of a Teflon tape (except of the side facing the PMT and area facing the source) serving as a diffuse light reflector and as additional thermal insulation to prevent cracking of the crystal due to any possible sudden changes in temperature of the sample holder. Then it was secured in the sample holder using a few additional pieces of a Teflon tape.
@@ 3.1 Light yield over 5–300 K temperature interval
The mean of the fit provided a relative measurement of the light yield. The peak position of the ’s was found to be 30,650 250 ADUs with a standard deviation of 1,800 300 ADUs while ’s peak position was determined to be 65,400 100 ADUs with a standard deviation of 3,240 40 ADUs. The uncertainties provided for light yield measurements includes only statistical uncertainties associated with the best fit parameters.
@@ 3.2 Pulse-shape analysis
Initially, events with a noisy baseline, pileups, and badly determined staring point were discarded from the collected dataset. Then, the different peak position on the detector energy scale for events induced by and irradiation was used to preliminary distinguish and select events of each type. See, for instance, the corresponding peak positions on the energy spectrum shown in figure 3. Events within two sigma around mean value of the corresponding peaks were selected for the further analysis. Only events with energy corresponding to a 662 keV full absorption peak were further considered as events in the pulse-shape analysis. Next, the mean time and prompt-to-total collected charge ratio were calculated for each event, and were used as additional discrimination and selection parameters to improve quality of the selected “pure” event of each type. These methods are well described in [31], where authors adopted those techniques to discriminate signals induced by particles from those of quanta in measurements with a CHC crystal.
Once “pure” events of each irradiation type were identified, the corresponding average waveforms were calculated. Then, the average waveform was fitted with a triple exponential function described by eq. 3.1, where , as previously shown in [13]. To fit the average ’s waveform was used a double exponential function provided by eq. 3.2, where . In both eq. 3.1 and eq. 3.2, represents the time duration of the average pulse, is the decay time-constant, and is the normalized relative intensity. The best fit of both types average waveforms normalized to their area, see figure 5, indicates that the average ’s waveform lasts longer than the average ’s waveform.
@@ 4 Discussion and conclusion
The intensity of the scintillating light emission under particles irradiation is rising when the CZC crystal is cooling down from RT to about 135 K reaching 19,700 1,700 photons/MeV, and then the amount of the emitted light decreases to about 10,600 900 photons/MeV at 75 K. While in the temperature interval from 75 K down to 5 K, it remains constant within uncertainties. It should be emphasized, that the present measurements confirms unusual behavior of the scintillating light intensity observed with the CZC crystal in [13], a decrease in the luminescence intensity as the temperature decreases from 125 K to 80 K, referred there as a “negative quenching”.